A generic salt, AB2, has a molar mass of 267 g/mol and a solubility of 4.10 g/L at 25 °C. What is the Ksp of this salt at 25 °C?
Answer #1

Ksp = [A+2][B-]^2

molar S = mass/MW = 4.1/267 = 0.015355

Ksp = (0.015355)(0.015355*2)^2

Ksp = 0.00001448136

Ksp = 1.48*10^-5